2. The owner of a factory that sells a particular bottled fruit juice claims that the average capacity of their product is 250 ml. To test the claim, a consumer group gets a sample of 100 such bottles, calculates the capacity of each bottle, and then finds the mean capacity to be 248 ml. The standard deviation is 5 ml. Is the claim true?

1
Expert's answer
2022-01-27T08:33:19-0500

Here,

"\overline{x}=248"

"n=100"

"\alpha =0.5"

"\:s=5"

Hypothesis tested are,

"H_o:\mu=250"

"vs"

"H_1:\mu\:\not=" "250"

The test statistic is,

"t=\frac{\overline{x}-\mu }{\frac{s}{\sqrt{n}}}=\frac{248-250}{\frac{5}{\sqrt{100}}}={-2\over{1\over2}}=-2\times2=-4."

Therefore,

"t=-4"

and

"\left|t\right|=4"

The critical table value, "t_{{\alpha\over2},n-1}" with "\alpha=0.05" with "df=n-1=99" is "t_{0.025,99}= 1.984217".

The null hypothesis is rejected if "|t|\gt t_{0.025,99}"

Here "\left|t\right|=4>t_{0.025,99}=1.984217"

Therefore, we reject the null hypothesis at "5\%"% level of significance and conclude that there is no sufficient evidence to show that the average capacity of the product is 250 ml.